3) Find the complete integral of p^3+q^3=27Z.


1
Expert's answer
2021-09-29T04:38:07-0400

"\u2212\n27\nz\n+\nq\n3\n+\np\n3"

"(\n\u2212\n27\nz\n+\nq\n3\n+\np\n3\n)\nx\n+\nC"

"\int{\left(p^{3} + q - 27 z + 3\right)d p} = \frac{p \left(p^{3} + 4 q - 108 z + 12\right)}{4}"

answer;

"\int{\left(p^{3} + q - 27 z + 3\right)d p}=\frac{p \left(p^{3} + 4 q - 108 z + 12\right)}{4}+C"